OpenStudy (anonymous):

Which value is an extraneous solution to the equation, |x| = -5x +14 -6 4 6 7 I don't understand this question because I believe three of those numbers are extraneous solutions. Only 4 fits

OpenStudy (solomonzelman):

You will have to different solutions. |x| can mean -x or x so you would have the following equations where in each you solve for x: -x = -5x +14 & x = -5x +14 (Do you see the difference?) (Need more help?)

OpenStudy (anonymous):

yes please

OpenStudy (solomonzelman):

Lets go with -x = -5x +14 Add 5x to both sides, can you do that?

OpenStudy (anonymous):

-5 = -5(5) + 14 -5 = -25 + 14 -5 = -11

OpenStudy (solomonzelman):

why are you doing that? Don't plug in numbers add 5x to both sides -x +5x= -5x +5x +14 right?

OpenStudy (anonymous):

then you would get 14= 4x ?

OpenStudy (solomonzelman):

Yes and if you plug that into your original equation, does that work or not?

OpenStudy (solomonzelman):

I mean x=14/4

OpenStudy (solomonzelman):

Plug it in and tell me what you think

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

So it is not extraneous solution, b/c it does not work sets work the second choice x = -5x +14

OpenStudy (anonymous):

you would get a decimal

OpenStudy (solomonzelman):

yes, and in any case w/ a second choice you would also, just by looking at it I know that none of your choices are right!

OpenStudy (solomonzelman):

x = -5x +14 add 5x to both sides, what do you get for x?

OpenStudy (anonymous):

6x= 14 14/6 +2.3333

OpenStudy (anonymous):

I don't understand where this is going. How do I plug in the answers above.

OpenStudy (solomonzelman):

|x| = -5x +14 -6 4 6 7 |-6| = -5(-6) +14 6= 30 +14 |4| = -5(4) +14 4=-6 |6| = -5(6) +14 6=-16 |7| = -5(7) +14 7= -35+14 7=-21 None of them work, right?

OpenStudy (solomonzelman):

@hottiewithabody, if you plug 14/6 you will see that it works perfectly.